# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ Abstract
Externally controlled single-arm trials are critical to assess treatment efficacy across therapeutic indications for which randomized controlled trials are not feasible. A closely-related research design, the unanchored indirect treatment comparison, is often required for disconnected treatment networks in health technology assessment. We present a unified causal inference framework for both research designs. We develop a novel estimator that augments a popular weighting approach based on entropy balancing – matching-adjusted indirect comparison (MAIC) – by fitting a model for the conditional outcome expectation. The predictions of the outcome model are combined with the entropy balancing MAIC weights. While the standard MAIC estimator is singly robust where the outcome model is non-linear, our augmented MAIC approach is doubly robust, providing increased robustness against model misspecification. This is demonstrated in a simulation study with binary outcomes and a logistic outcome model, where the augmented estimator demonstrates its doubly robust property, while exhibiting higher precision than all non-augmented weighting estimators and near-identical precision to G-computation. We describe the extension of our estimator to the setting with unavailable individual participant data for the external control, illustrating it through an applied example. Our findings reinforce the understanding that entropy balancing-based approaches have desirable properties compared to standard “modeling” approaches to weighting, but should be augmented to improve protection against bias and guarantee double robustness.
@@ 1 Background
In pharmaceutical research, randomized controlled trials (RCTs) are the gold standard for evaluating the efficacy and safety of new treatments, due to their high internal validity. The random allocation of subjects to either an experimental therapy or a concurrent control minimizes the risk of confounding, balancing in expectation the distribution of prognostic factors over the treatment arms. Nevertheless, conducting an RCT is not possible in certain settings. For instance:
@@
  • Where recruitment to properly-powered RCTs is unfeasible or impractical due to small populations and a lack of trial-eligible patients, e.g., rare diseases with orphan designation or biomarker-specific precision oncology indications;1
@@
  • For life-threatening and severely debilitating conditions with high unmet medical need, due to non-existent or inadequate standard of care, e.g., “last-line of therapy” indications in late-stage hematological and solid tumor oncology;2
@@
  • Where the enrollment of patients to a placebo or “no treatment” control arm is unethical, e.g., withholding a therapy with proven efficacy in adults from a control group of children in pediatric trials.3
@@
Regulatory agencies such as the Food and Drug Administration (FDA) and the European Medicines Agency (EMA) emphasize that properly conducted RCTs provide the highest evidentiary standard for the approval of new drugs.4, 5 Nevertheless, regulators recognize that alternative research designs may be required in special circumstances. One of such designs is the externally controlled single-arm trial (SAT), where the control group is fully derived from external data such as prior clinical trials or secondary real-world data (RWD) sources, e.g., electronic health records, disease registries or medical claims.2, 6, 7, 8
Marketing authorization applications featuring externally controlled SATs continue to rise, especially under the accelerated approval pathway from the FDA and the conditional approval pathway from the EMA.8, 9 In 2023 the FDA issued a draft guidance document for externally controlled trials,10 and in 2024 the EMA finalized a reflection paper on single-arm trials.11 Regulators have granted approvals to new therapeutics based on externally controlled SATs, especially in rare diseases, oncology and hematology, with approval being more likely for conditions with a highly predictable natural history, precisely measurable endpoints, and where a large effect size is anticipated.1, 2, 8, 9, 12, 13
For health technology assessment (HTA), head-to-head RCT evidence also remains the gold standard for decision-making.14, 15, 16 Nevertheless, as regulatory authorities increasingly approve pharmaceuticals based on externally controlled SATs, the reliance of payers on such research designs has grown.9, 17 HTA bodies are developing recommendations for externally controlled SATs in reimbursement submissions,18, 19 with acceptability being greater in oncology and also influenced by factors such as high unmet need and the rarity of the disease.9, 17, 20
Additionally, HTA requires comparing the clinical effectiveness and cost-effectiveness of new health technologies against all existing alternatives on the market.21 The scope of assessments often depends on the policy question and is not always driven by the available data.22 A single RCT cannot typically have all the treatment arms that are desired for HTA, particularly given the multiplicity of national and regional stakeholders, and the variations in clinical practice across jurisdictions, which may themselves arise due to historical variations in reimbursement practice.23 In the absence of direct RCT comparisons versus all candidate comparators, indirect treatment comparisons (ITCs) across studies are required.24
@@
HTA decision-makers have expressed a clear preference for anchored ITCs of randomized trials.25, 26, 27 These respect study randomization by using a common control arm to contrast relative treatment effects across trials.28 Nevertheless, even for indications where RCTs are regularly conducted, at times it is not possible to find a compatible control arm with which to “anchor” the analysis. This is especially the case in rapidly evolving therapeutic areas with multiple novel treatments entering the market, a changing comparator landscape and no single accepted standard of care internationally.29, 30 In these scenarios, unanchored ITCs based on disconnected treatment networks may be required29, 30 and a recent literature review determined that unanchored ITCs are in fact far more common than anchored ITCs. 31
Unanchored ITCs contrast mean treatment-specific absolute outcomes across studies, thereby relying on more restrictive assumptions than anchored ITCs.25, 28 In essence, unanchored ITCs are externally controlled SATs with two specific characteristics. Firstly, the external control is often a historical trial that has been sponsored by a competitor. Secondly, due to privacy and confidentiality reasons, there is limited access to subject-level data. While individual participant data (IPD) are available for the SAT, only published aggregate-level data (AD) are available for the external control.25, 28
@@
The absence of randomization compromises the validity of externally controlled SATs. Various statistical methods have been proposed to adjust for imbalances in baseline covariates between the SAT and the external control.32 These can potentially mitigate confounding bias and account for the additional variability induced by covariate differences. The most widely-used techniques are propensity score-based weighting approaches, which explicitly model the conditional probability of SAT participation – a propensity score – as a function of baseline covariates. Typically, a logistic regression is applied, such that the logit of the propensity score is assumed to vary linearly with the covariates, and this is estimated via maximum-likelihood.33, 34, 35, 36
Specifically within the context of unanchored ITCs, an alternative weighting method based on entropy balancing37 called matching-adjusted indirect comparison (MAIC) is more popular.38, 39, 40 This views covariate balance as a convex optimization problem, estimating weights that directly enforce balance in covariate moments between the SAT and the external control, without explicitly modeling the conditional probability of SAT participation. MAIC is attractive for ITCs due to its applicability in the “IPD-AD” situation, where there is limited access to IPD for the external control. Moreover, entropy balancing techniques such as MAIC are thought to be generally more stable, more precise, and more robust to model misspecification relative to the standard “modeling” approaches to weighting, even in “IPD-IPD” scenarios.40, 41, 42, 43
@@
So-called “G-computation” or “model-based standardization” methods have also been developed for the IPD-IPD33, 34, 44, 45 and IPD-AD settings.46 For the latter, they are also referred to as “simulated treatment comparison” (STC).47 These methods are based on estimating a model for the conditional outcome expectation given baseline covariates (the “outcome model”), and averaging model-based outcome predictions over the target covariate distribution. G-computation approaches gain statistical precision and efficiency with respect to weighting, particularly where poor covariate overlap leads to extreme weights and large reductions in effective sample size.46 That being said, the improved performance relies on the extrapolation of a correctly specified outcome model and G-computation methods can be prone to bias and undercoverage, even under relatively minor model misspecification.48
@@
Weighting and G-computation techniques make different modeling assumptions, but both are generally “singly robust”. The former, in most cases, depends on correct specification of the propensity score model for consistent – and therefore asymptotically unbiased – estimation. The latter depends on correct specification of the model for the conditional outcome expectation. Decision-makers have expressed a preference for “doubly robust” estimation approaches that can consistently estimate the treatment effect as long as either the propensity score model or the outcome model is correct, but not necessarily both.25, 27, 28, 49 These methods should reduce the risk of bias by offering two opportunities for correct model specification. Despite this, doubly robust methods have rarely been applied in the analysis of externally controlled SATs.50 To the best of our knowledge, they have never been applied to unanchored ITCs, despite their development being recommended by HTA agencies.25, 28
@@
One reason why doubly robust methods do not feature prominently for unanchored ITCs might be a misunderstanding that MAIC, the most popular approach for ITCs,31 is always doubly robust. To be clear, MAIC does enable consistent estimation when an implicit propensity score model is misspecified, but only if the true outcome model is linear with respect to the covariate functions that are balanced. As such, entropy balancing is said to be “linearly doubly robust”,40, 42, 43 with Zhao and Percival (2017) stating that it is “doubly robust with respect to linear outcome regression and logistic propensity score regression”.43 In practice, it is rarely plausible that outcomes vary linearly with the covariates (e.g., due to non-linear link functions or outcomes that depend on non-linear transformations of the covariates). Doubly robust methods for ITCs that are not necessarily restricted to linear outcome models are yet to be developed, with Josey et al (2021) recently identifying this as a research priority.40
The objective of this paper is to clarify existing approaches for doubly robust estimation in the analysis of externally controlled SATs and propose a doubly robust augmented MAIC estimator, specifically tailored to unanchored ITCs. In Section 2, we introduce and motivate the target estimand. In Section 3, we outline the available singly and doubly robust estimators and our proposed doubly robust estimator for unanchored ITCs, explaining the rationale behind the latter being doubly robust. In Section 4, we empirically evaluate the statistical performance of the approaches in a simulation study, comparing doubly robust estimators against the singly robust estimators and other augmented estimators that have also been described as doubly robust, such as G-computation with an outcome model fitted using weighted maximum-likelihood estimation.51, 52 The simulations demonstrate proof-of-principle with binary outcomes, a logistic outcome model and the marginal log-odds ratio as the target summary effect measure, where entropy balancing approaches do not have the doubly robust property. In Section 5, we illustrate the application of the methods in an example analysis. Lastly, we discuss our findings and conclude the paper in Section 6.
@@ 2 Estimands
Prior to introducing the statistical methodology, we define the different estimands that can be targeted by externally controlled SATs. An estimand is a precise definition of the treatment effect, which should align with the clinical question of interest, the research design and the analytical approach. Within the regulatory environment, the use of estimands has been stimulated by the publication of the International Council of Harmonisation E9 (R1) Addendum, adopted by the FDA and the EMA. While E9 (R1) was originally proposed for registrational RCTs, its principles are applicable to externally controlled SATs.53 According to E9 (R1), an estimand encompasses five attributes: target population, treatment(s), endpoint (outcome), summary effect measure and the strategy for handling intercurrent events.
We will focus specifically on the “target population” and “summary effect measure” components when defining the estimands, which are:
@@
  • The average treatment effect in the treated (ATT); that is, among those participating in the SAT;
  • The average treatment effect in the control (ATC); that is, among those in the external control group.
The difference between these summary effect measures is driven by them targeting different (sub) populations or applying to different “analysis sets”. Having assumed that the SAT and external control analysis sets are random samples of their underlying target populations, we will generally make no further distinction between sample-level and population-level estimands.
We adopt the potential outcomes notation to define the estimands more formally. Let $Y^{t}$ represent the potential outcome that would have been observed for a subject under intervention $T=t$, with $t\in\{0,1\}$, where $T=1$ denotes an active intervention under investigation in the SAT ($S=1$), and $T=0$ denotes a pertinent control for the subjects in the external data source ($S=0$). Two potential outcomes, $(Y^{1},Y^{0})$, are defined for every subject, regardless of whether they are members of the SAT under $T=1$ or of the external control under $T=0$. Assuming no loss to follow up, one of the potential outcomes is observed for a given subject and the other is the “counterfactual” which would be realized under a different treatment than that actually assigned.
The ATE is defined as:
@@
on the additive scale imposed by link function $g(\cdot)$, which transforms the potential outcome means into the plus/minus infinity range. Each expectation, $\textrm{E}(\cdot)$, is taken over the distribution of potential outcomes in the combined SAT and external control population. For a binary outcome, suitable link functions could be the identity, log or logit, to produce a risk difference, log relative risk or log-odds ratio as the summary effect measure. The ATT is defined as:
@@
such that expectations are taken over the distribution of potential outcomes in the SAT (sub) population. The ATC is defined as:
@@
such that expectations are taken over the distribution of potential outcomes in the external control (sub) population.
Within RCTs, the ATE, ATT and ATC are identical in expectation. However, they generally differ in externally controlled SATs, and will almost invariably do so where there is (conditional) treatment effect heterogeneity by the covariates, i.e., effect measure modification. We view the ATE target population, defined by pooling the SAT and the external control, as somewhat ambiguous in this context. As such, the target estimand in an externally controlled SAT is often either the ATT or the ATC.
The ATT is typically the primary estimand for drug approval purposes in the regulatory environment. It is consistent with emulating a randomized comparison in the pivotal trial population, with the external control mimicking the internal comparator arm of a registrational clinical trial. The ATT is also compatible with the mean absolute outcome that is targeted by the SAT, $\textrm{E}(Y^{1}\mid S=1)$, preserving the original SAT results. Nevertheless, SAT populations are often highly selected and may lack representativeness with respect to “real-world” patient populations. As such, the ATT may be less appealing where generalizability to routine clinical practice is a priority, as is the case in HTA decision-making.
The ATC is potentially more desirable for external validity. External controls based on natural history studies and RWD have broad inclusion criteria and populations that are relatively heterogeneous in clinical and demographic characteristics. However, external controls are not always representative of routine clinical practice. Historical controls from past clinical trials will not reflect the current standard of care due to the passage of time. Moreover, RWD-derived external controls are often based on a single country and not necessarily transferable to the relevant jurisdiction for decision-making.
Sample size considerations and challenges in statistical estimation may also play a part in the estimand choice. Both SATs and external controls often have low sample sizes, particularly in rare diseases. Covariate adjustment may reduce effective sample sizes further. The ATT or ATC may not be reliably estimated where sample size is low in the external control or the SAT, respectively. Consider weighting, where the estimand impacts the definition of the weights. Targeting the ATT implies preserving the original SAT, re-weighting and reducing the effective sample size of the external control. Conversely, targeting the ATC implies preserving the original external control, re-weighting and reducing the effective sample size of the SAT.
Finally, where IPD are available for the SAT but only published AD are available for the external control, as is typically the case in unanchored ITCs, the target estimand is often the ATC by necessity.25, 28, 32 Throughout the rest of the manuscript, we shall assume that there is unlimited access to subject-level data but that the ATC is the primary target of estimation. Our methodological approaches and findings are also applicable to the ATT, or where subject-level data are unavailable for the external control, with some caveats that will be discussed in Section 3.

### v2

@@ Abstract
Externally controlled single-arm trials are critical to assess treatment efficacy across therapeutic indications for which randomized controlled trials are not feasible. A closely-related research design, the unanchored indirect treatment comparison, is often required for disconnected treatment networks in health technology assessment. We present a unified causal inference framework for both research designs. We develop an estimator that augments a popular weighting approach based on entropy balancing – matching-adjusted indirect comparison (MAIC) – by fitting a model for the conditional outcome expectation. The predictions of the outcome model are combined with the entropy balancing MAIC weights. While the standard MAIC estimator is singly robust where the outcome model is non-linear, our augmented MAIC approach is doubly robust, providing increased robustness against model misspecification. This is demonstrated in a simulation study with binary outcomes and a logistic outcome model, where the augmented estimator demonstrates its doubly robust property, while exhibiting higher precision than all non-augmented weighting estimators and near-identical precision to G-computation. We describe the extension of our estimator to the setting with unavailable individual participant data for the external control, illustrating it through an applied example. Our findings reinforce the understanding that entropy balancing-based approaches have desirable properties compared to standard “modeling” approaches to weighting, but should be augmented to improve protection against bias and guarantee double robustness.
@@
Highlights
What is already known?
  • Externally controlled single-arm trials are important for assessing treatment efficacy when randomized controlled trials are not feasible.
  • Unanchored indirect treatment comparisons, often using the entropy balancing or matching-adjusted indirect comparison (MAIC) approach, can provide treatment effect estimates for disconnected networks in health technology assessment.
What is new?
  • A unified causal inference framework for both externally controlled single-arm trials and unanchored indirect treatment comparisons is presented.
  • A doubly robust estimator that augments the MAIC approach by combining outcome model predictions with entropy balancing weights is proposed.
  • The proposed estimator provides increased protection against model misspecification, as demonstrated in a simulation study with binary outcomes.
  • The proposed estimator is extended to the setting with unavailable individual participant data for the external control and illustrated through an applied example.
Potential impact for RSM readers outside the authors’ field
  • Doubly robust augmented weighting estimators, particularly using entropy balancing weights, offer increased robustness to model misspecification than conventional non-augmented approaches for confounding adjustment.
  • Entropy balancing-based methods have desirable properties compared to standard “modeling” approaches to weighting but should be augmented to improve protection against bias and guarantee double robustness.
@@ 1 Background
In pharmaceutical research, randomized controlled trials (RCTs) are the gold standard for evaluating treatment efficacy and safety due to their high internal validity. Random allocation minimizes confounding by balancing, in expectation, prognostic factors across treatment arms. Nevertheless, RCTs are not always feasible, for instance:
@@
  • Where recruitment is impractical due to small populations, e.g., rare diseases or biomarker-specific precision oncology;jahanshahi2021use
@@
  • For life-threatening conditions with high unmet need and inadequate standard of care, e.g., “last-line of therapy” indications in late-stage hematological and solid tumor oncology;mishra2022external
@@
  • Where placebo controls are unethical, e.g., withholding a therapy with proven efficacy in adults from a control group of children in pediatric trials.horton2021real
@@
Regulatory agencies such as the Food and Drug Administration (FDA) and the European Medicines Agency (EMA) emphasize that RCTs provide the highest evidentiary standard.beaulieu2020examining, flynn2022marketing However, regulators recognize that alternative designs may be required in special circumstances. One such design is the externally controlled single-arm trial (SAT), where the control group is fully derived from external data such as prior clinical trials or secondary real-world data (RWD) sources.mishra2022external, zou2024next, curtis2023regulatory, gray2020framework Marketing authorization applications featuring externally controlled SATs continue to rise, especially under accelerated approval pathways.gray2020framework, sola2023effectively In 2023 the FDA issued draft guidance for externally controlled trials,FDA2023guidance and in 2024 the EMA finalized a reflection paper on SATs.EMA2024reflection Approvals have been granted based on such designs, particularly for conditions with predictable natural history, precisely measurable endpoints, and anticipated large effect sizes.jahanshahi2021use, mishra2022external, gray2020framework, sola2023effectively, bakker2023contribution, goring2019characteristics
For health technology assessment (HTA), head-to-head RCT evidence also remains the gold standard.NICE2022manual, CADTH2024procedures, IQWiG2023methods Nevertheless, as regulators increasingly approve pharmaceuticals based on externally controlled SATs, payer reliance on such designs has grown.sola2023effectively, patel2021use HTA bodies are developing recommendations for externally controlled SATs,NICE2022framework, IQWiG2020guidance with acceptability influenced by unmet need and disease rarity.sola2023effectively, patel2021use, jaksa2022comparison HTA also requires comparing new technologies against all existing alternatives.paul2001fourth The scope of assessments often depends on the policy question rather than available dataEUnetHTA2021joint and a single RCT cannot typically include all comparators desired for HTA, given the multiplicity of stakeholders and variations in clinical practice across jurisdictions.vreman2020differences In the absence of direct RCT comparisons versus all candidate comparators, indirect treatment comparisons (ITCs) are required.sutton2008use
@@
HTA decision-makers prefer anchored ITCs of randomized trials,phillippo2016nice, EUnetHTA2022methods, EUnetHTA2022practical which respect randomization by using a common control arm to contrast relative treatment effects.phillippo2018methods However, compatible control arms with which to “anchor” the analysis are not always available, especially in rapidly evolving areas with multiple novel treatments and no single accepted standard of care.goring2016disconnected, stevens2018review In these scenarios, unanchored ITCs based on disconnected networks may be required,goring2016disconnected, stevens2018review and recent reviews determined that unanchored ITCs are in fact more common than anchored ITCs.serret2023methodological, truong2023population Unanchored ITCs contrast mean treatment-specific absolute outcomes across studies, relying on more restrictive assumptions than anchored ITCs.phillippo2016nice, phillippo2018methods In essence, they are externally controlled SATs where the external control is often a competitor’s historical trial with limited data access: individual participant data (IPD) are available for the SAT, but only published aggregate-level data (AD) for the external control.phillippo2016nice, phillippo2018methods
@@
The absence of randomization compromises the validity of externally controlled SATs. Various statistical methods have been proposed to adjust for imbalances in baseline covariates lambert2023enriching and these can potentially mitigate confounding bias and account for the additional variability induced by covariate differences. The most widely-used methods are propensity score-based weighting approaches, typically using logistic regression estimated via maximum-likelihood.ren2023comparing, loiseau2022external, lunceford2004stratification, austin2016variance For unanchored ITCs, matching-adjusted indirect comparison (MAIC) based on entropy balancinghainmueller2012entropy is more popular.signorovitch2010comparative, cheng2020statistical, josey2021transporting MAIC views covariate balance as a convex optimization problem, estimating weights that directly enforce balance without explicitly modeling the conditional probability of SAT participation. MAIC is attractive for ITCs due to its applicability in IPD-AD situations and is thought to be more stable, precise, and robust to model misspecification than the standard propensity score-based weighting approaches, even in “IPD-IPD” scenarios.josey2021transporting, amusa2019examination, cheng2023double, zhao2017entropy
@@
So-called “G-computation” or “model-based standardization” methods have also been developed for the IPD-IPDloiseau2022external, wang2017g, robins1986new, keil2014parametric and IPD-AD settings.remiro2022parametric For the latter, they are also referred to as “simulated treatment comparison” (STC).ishak2015simulation These methods estimate a model for the conditional outcome expectation and average predictions over the target covariate distribution. G-computation exhibits increased precision relative to weighting, particularly when overlap is poor,remiro2022parametric but relies on model-based extrapolation and can be prone to bias under model misspecification.vo2023cautionary
@@
Weighting and G-computation are generally “singly robust”: weighting, in most cases, depends on correct propensity score model specification; G-computation on correct outcome model specification. Decision-makers have expressed a preference for “doubly robust” estimation approaches that can consistently estimate the treatment effect as long as either the propensity score model or the outcome model is correct, but not necessarily both.phillippo2016nice, EUnetHTA2022practical, phillippo2018methods, vanier2024rapid These methods should reduce the risk of bias by offering two opportunities for correct model specification. Despite this, doubly robust methods have rarely been applied to externally controlled SATssiu2024framework and, to our knowledge, never to unanchored ITCs, despite recommendations from HTA agencies.phillippo2016nice, phillippo2018methods
@@
One barrier may be a misunderstanding that MAIC is always doubly robust. MAIC enables consistent estimation when an implicit propensity score model is misspecified, but only if the true outcome model is linear with respect to balanced covariate functions – termed “linearly doubly robust”.josey2021transporting, cheng2023double, zhao2017entropy In practice, outcomes rarely vary linearly with covariates. Doubly robust methods for ITCs that are not necessarily restricted to linear outcome models are yet to be developed, with Josey et al (2021) recently identifying this as a research priority.josey2021transporting
This paper clarifies existing approaches for doubly robust estimation in externally controlled SATs and proposes a doubly robust augmented MAIC estimator for unanchored ITCs. Section 2 introduces the target estimand. Section 3 outlines available estimators and our proposed approach. Section 4 presents simulation study results comparing our proposed doubly robust augmented estimators against existing singly robust and other augmented estimators. Lastly, we illustrate the application of the methods in an example analysis in Section 5, and conclude in Section 6.
@@ 2 Estimands
We begin by defining the estimands that can be targeted by externally controlled SATs. An estimand is a precise definition of the treatment effect, which should align with the clinical question of interest, the research design and the analytical approach. The International Council of Harmonisation E9 (R1) Addendum, adopted by the FDA and EMA, specifies five estimand attributes: population, treatment(s), endpoint, summary effect measure, and strategies for intercurrent events.polito2024applying We focus on the “population” and “summary effect measure” when defining the estimands, which are:
@@
  • The average treatment effect in the treated (ATT), among those participating in the SAT; and
  • The average treatment effect in the control (ATC), among the external control group.
The difference between these summary effect measures is driven by them targeting different (sub) populations or applying to different “analysis sets”. Having assumed the SAT and external control are random samples of their underlying target populations, we make no further distinction between sample-level and population-level estimands.
Using potential outcomes notation, let $Y^{t}$ represent the outcome under intervention $T=t$, with $t\in\{0,1\}$, where $T=1$ denotes the SAT intervention (data source $S=1$) and $T=0$ the external control (data source $S=0$). Two potential outcomes, $(Y^{1},Y^{0})$, exist for every subject; one is observed, the other counterfactual. The ATE is:
@@
where the link function $g(\cdot)$ transforms potential outcome means into the plus/minus infinity range, and expectations are over the distribution of potential outcomes in the combined SAT and external control population. For binary outcomes, suitable links include identity, log or logit, to produce a risk difference, log relative risk or log-odds ratio, respectively, as the summary effect measure. The ATT is:
@@
with expectations taken over the SAT (sub) population. The ATC is:
@@
with expectations over the external control (sub) population.
Within RCTs, the ATE, ATT and ATC are identical in expectation. However, they generally differ in externally controlled SATs, and will almost invariably do so where there is treatment effect heterogeneity by the covariates, i.e., effect measure modification. We view the ATE target population, defined by pooling the SAT and the external control, as somewhat ambiguous in this context. As such, the target estimand in an externally controlled SAT is often either the ATT or the ATC.
The ATT is typically the primary estimand for regulatory drug approval, consistent with emulating a randomized comparison in the pivotal trial population, with the external control mimicking the internal comparator arm of a registrational clinical trial. The ATT is also compatible with the mean absolute outcome that is targeted by the SAT, $\textrm{E}(Y^{1}\mid S=1)$, preserving the original SAT results. Nevertheless, SAT populations are often highly selected and may lack “real-world” representativeness, making the ATT potentially less appealing for HTA, where generalizability to routine clinical practice is a priority.
The ATC can be more desirable for external validity, as natural history and RWD-based external controls have relatively broad inclusion criteria and heterogeneous target populations. However, external controls based on historical trials will not reflect the current standard of care and RWD-derived controls are often country-specific, not necessarily transferable to the relevant jurisdiction for decision-making. Sample size considerations also influence the estimand choice. Both SATs and external controls often have low sample sizes and covariate adjustment may reduce effective sample sizes further. Consider weighting, where the estimand impacts the definition of the weights. Targeting the ATT implies preserving the original SAT, re-weighting and reducing the effective sample size of the external control. Conversely, targeting the ATC implies the reverse.
Finally, where IPD are available for the SAT but only AD for the external control, as in unanchored ITCs, the ATC is often targeted by necessity.phillippo2016nice, phillippo2018methods, lambert2023enriching Throughout this manuscript, we assume unlimited subject-level data access but target the ATC. Our methodological approaches are also applicable where subject-level data are unavailable for the external control, and to target the ATT instead of the ATC, with caveats discussed in Section 3.8 and the Supplementary Material, respectively.